OpenStudy (anonymous):

Gives medals ! Write the equation of the line that passes through (–2, 6) and (2, 14) in slope-intercept form.

OpenStudy (anonymous):

@jigglypuff314 @Michele_Laino

jigglypuff314 (jigglypuff314):

slope intercept form is \( y=mx+b\) where m = slope and b = y-intercept

jigglypuff314 (jigglypuff314):

so from the given points, we have to first find the slope do you know how to do that? :)

OpenStudy (anonymous):

No, i don't quiet understand any of this. These are the multiple choice . y = 2x – 2 y = 2x + 10 y = 0.5x + 7 y = 0.5x – 4

jigglypuff314 (jigglypuff314):

slope = (y2 - y1) / (x2 - x1) when given two points (x1, y1) and (x2, y2) have you ever seen a rule like this before?

OpenStudy (anonymous):

Yes I have. hold on I will find the slope.

OpenStudy (anonymous):

The slope is 2.

jigglypuff314 (jigglypuff314):

perfect! :) then we can plug that into our equation and get y = 2x + b so now all we have to do is find "b"

OpenStudy (anonymous):

and B is 10

OpenStudy (anonymous):

So is would be answer B!?

jigglypuff314 (jigglypuff314):

that's right! :)

OpenStudy (anonymous):

Thanks so much you're a great help!

jigglypuff314 (jigglypuff314):

I'm glad I could help ^_^

jigglypuff314 (jigglypuff314):

sh shh shhh delete that you could get banned if the mods found out you got help on a test ;)

OpenStudy (anonymous):

Got it (;